OpenStudy (anonymous):

name the slope and y intercept of y = -3/5x + 1

OpenStudy (anonymous):

Slope: -3/5, y-int = 1

OpenStudy (amistre64):

hmmm, if there was an equation that utilized the slope and y intercept we could easily get this

jimthompson5910 (jim_thompson5910):

y = -3/5x + 1 is in y = mx+b form where m is the slope and (0, b) is the y-intercept Since m = -3/5, the slope is -3/5 Because b = 1, the y-intercept is (0, 1)

OpenStudy (saifoo.khan):